# Knowledge Enhanced Conditional Imputation for Healthcare Time-series

### v3

@@ 4.3 Non-Uniform Masking Strategy
  • $\boldsymbol{R_{\text{factor}}}(d|U,I)$ is the regulatory factor for feature $d$, conditioned by $U$ and $I$.
  • $\boldsymbol{P_{\text{dist}}}(d)$ represents the probability distribution of missingness for feature $d$.
The overall masking proportion is then adjusted to ensure consistency with the uniform masking rate $U$, while retaining the non-uniform characteristics of the individual features. The pseudocode for this algorithm is as follows:
Algorithm 1 Non-Uniform Masking Strategy
Input: Dataset $\boldsymbol{X}$ with $D$ features, Uniform masking rate $U$, weighting parameter $I$
@@
for feature $d$ in $D$ do
     $\boldsymbol{P_{\text{dist}}}(d)\leftarrow\boldsymbol{x^{d}}$
@@
     $\boldsymbol{P_{\text{nu}}}(d)\leftarrow R_{\text{factor}}(d)\times\boldsymbol{P_{\text{dist}}}(d)$
@@
$U\leftarrow f(\boldsymbol{P_{\text{nu}}}(d),\boldsymbol{x^{d}})$
@@
5 Experimental Evaluation
We conducted a number of experiments to evaluate the proposed methodologies and their effect on CSAI’s performance. First, we compared CSAI’s overall imputation and classification performance against five state-of-the-art baselines using four real-world healthcare datasets. The comparison includes BRITS, BRITS with GRU (BRITS-GRU), GRU-D, V-RIN, and M-RNN. We included BRITS-GRU alongside the original BRITS, which uses LSTM cells, to ensure a fair comparison, as the other baselines employ GRU cells. This addition allows for a more comprehensive and balanced evaluation across similar architectures. Although a comparison with $E^{2}$ would have provided valuable insight, particularly given its lack of quantitative comparison with BRITS, the publicly available implementation of $E^{2}$ is incompatible with our current setup. To ensure reproducibility of our results, all data preprocessing scripts, model implementations, and hyperparameter search configurations are released.
We then perform a two-fold ablation study using the openly-available Physionet dataset to evaluate the merit of the proposed non-uniform masking strategy. First, our study establishes the effect of non-uniform masking across different data partitions, namely the training, validation and test sets. The aim is to establish whether this approach indeed improves a model’s ability to generalise and handle missing data in a more realistic and robust manner. Finally, we evaluate the effect of the weighting parameter, which controls the degree of emphasis placed on different features in the non-uniform masking strategy. This parameter allows for fine-tuning of the feature representation during model training, directly influencing how the model handles missing data. This set of experiments evaluates the effect on CSAI’s imputation and classification performance under varied the weighting parameter values.
5.1 Datasets
Each of the four datasets chosen for the experimental evaluation has a different data distribution, especially the MIMIC-III database, from which we extracted two different datasets to model different tasks. We reproduced the benchmarking for these public datasets, skipping steps that remove all-NAN samples to retain the data with its original missingness.
  1. 1.
    eICU [^1] https://physionet.org/content/eicu-crd/2.0/ [24] is sampled from the eICU Collaborative Research Database, a multi-center database with anonymised hospital patient records. eICU is publicly available after registration, including completion of a training course in research with human subjects and signing of a responsible data use agreement. We followed the only benchmarking extract available for eICU [25], yielding a dataset comprising 30,680 samples measuring 20 features, 3 of which were categorical.
  2. 2.
@@ 5.1 Datasets
    MIMIC-III [^2] https://physionet.org/content/mimiciii/1.4/ [26] is an extensive, freely-available database of over 40,000 critical care patients. To complement our MIMIC experiments with heterogeneous feature types and data dimensionality, we followed two benchmark papers [27, 28] to produce two datasets. The first comprises 89 (10 of which are categorical) variables for 14,188 samples while the second captures 59 numerical variables for 21,128 samples.
  3. 3.
@@
    PhysioNet Challenge 2012 dataset. The Predicting Mortality of ICU Patients: The PhysioNet/Computing in Cardiology Challenge 2012, a public medical benchmarking dataset provided by [29], contains records of 4000 48-hour ICU stays, allowing unbiased evaluations of different model performances. This benchmark comprises 35 numerical variables for 3,997 samples.
5.2 Implementation Details
We trained our models on an HPC node with NVIDIA A100 40GB running Ubuntu 20.04.6 LTS (Focal Fossa). We used Python: 3.8.16. All package details can be found in our repository. For each of the five datasets, we created three versions by further masking 5%, 10% and 20% cells in addition to the missingness already present in the original datasets. These masked cells have known ground truths and will form the basis of our comparison of imputation performance.
We used the Adam optimiser and set the number of RNN hidden units to 108 for all models. The batch size is 64 for PhysioNet and 128 for the other datasets. Each dataset was normalised to zero mean and unit variance. Randomly, we selected 10% of each dataset for validation and another 10% for testing, training the models on the remaining data. For the imputation task, we masked 10% of observations in each dataset to serve as the ground truth, which was used as validation data. A 5-fold cross-validation method was implemented to evaluate the models. Imputation and classification performance was measured using the Mean Absolute Error (MAE) and Area Under the Receiver-Operator Curve (AUC-ROC) respectively. We adopted a non-uniform masking strategy for the training set in all datasets. The code for CSAI is available on GitHub: https://github.com/LINGLONGQIAN/CSAI.
5.3 Experimental Results
[…]
Comparison with State-of-the-Art Models Our evaluation of CSAI against five state-of-the-art models using four diverse healthcare data sets is shown in Tables 1 and 2. Table 1 shows imputation performance, where CSAI consistently achieved the lowest MAE compared to all other models. This is particularly evident in lower masking ratios on both the eICU and mimic_59f datasets, suggesting that CSAI effectively uses its conditional knowledge embedding under limited data conditions to accurately impute missing values.
@@ 5.3 Experimental Results
Table 1: Imputation performance using 5%, 10% and 20% masking ratios on the four datasets. In each setup, the model with the highest MAE is highlighted in bold.
eICU (MAE) MIMIC_59f (MAE)
5 10 20 5 10 20
BVRIN 0.2416066624555 0.2425388767962 0.2521443794081 0.15457482983029 0.1381774096223 0.3369691066028
BRITS 0.1669909016914 0.1705338793299 0.1768130684914 0.1519486919982 0.1402299219611 0.3403944741631
BRITS_GRU 0.172319898527 0.1712448395078 0.1769143760391 0.1479284907740 0.1419320843095 0.3416945429830
GRUD 0.2227370102738 0.2256000191395 0.2309779583607 0.3044727329354 0.2870424025309 0.4866963332628
MRNN 0.4703575678949 0.4799816897156 0.5006539067069 0.305728740144 0.2834192980232 0.4719812224131
CSAI 0.1596681888527 0.1614887556508 0.1663724494172 0.1311938059021 0.1129106347724 0.3097617904199
PhysioNet (MAE) MIMIC_89f (MAE)
5 10 20 5 10 20
BVRIN 0.261627214279 0.2737156899792 0.2999655482813 0.2802916356976 0.2981212480367 0.3249882376070
BRITS 0.2563447937357 0.2676201849977 0.2872153778145 0.2846978765928 0.3063578035340 0.3298037475677
BRITS_GRU 0.2512872199774 0.2621551395257 0.2828820523984 0.2833476046520 0.3020169393225 0.325679723132
GRUD 0.4940628434272 0.4977919324802 0.509515400807 2.162001471223 2.226354328684 2.300878604115
MRNN 0.5467116967567 0.5564676306450 0.5722952103217 0.5077789122169 0.5202164923417 0.5340895067570
CSAI 0.2460182594695 0.2574714476838 0.2747586037341 0.2845187575872 0.2984590455884 0.3277731324925
Table 2 further demonstrates CSAI’s strong classification performance, achieving competitive AUC scores across various masking ratios in the eICU, PhysioNet, and MIMIC_59f datasets. However, CSAI slightly underperforms in the MIMIC_89f compared to BRITS-GRU. Notably, this dataset contains a high proportion of categorical variables, for which the architectural modifications of CSAI may not be as advantageous and which highlights the need for further tuning. Nevertheless, the robustness of the BRITS architecture further validates its selection as the CSAI backbone.
@@
Table 2: Classification performance using 5%, 10% and 20% masking ratios on the four datasets. In each setup, the model with the highest AUC is highlighted in bold.
eICU (AUC) MIMIC_59f (AUC)
5 10 20 5 10 20
BVRIN 0.8876632792674 0.8842207935257 0.8846083295889 0.8327985224797 0.8331225972961 0.8273344786621
BRITS 0.8866858256973 0.8852105575201 0.8856887012643 0.8282097938406 0.8277898698471 0.8241398505926
BRITS_GRU 0.8894446123008 0.8866442109252 0.8860531724480 0.8318754575867 0.830741815613 0.826876535038
GRUD 0.8649508677163 0.8645586279780 0.8580064406861 0.8276771341543 0.8263960525470 0.8230358085344
MRNN 0.8778555135764 0.8762569153075 0.8734187899382 0.8210628460196 0.8170771816893 0.8128184260363
CSAI 0.8895208094995 0.8897650378954 0.8879479677632 0.8352381478170 0.8336724124061 0.8310903756879
PhysioNet (AUC) MIMIC_89f (AUC)
5 10 20 5 10 20
BVRIN 0.8343250745783 0.8291631265561 0.8254836447961 0.8515838497027 0.8483391172226 0.8399032675067
BRITS 0.822140767045 0.811750067602 0.8218051204357 0.8518059085347 0.8499103429115 0.8410220558148
BRITS_GRU 0.8067688302396 0.8192578961191 0.8065060341956 0.85410308259 0.852057183334 0.84306840205
GRUD 0.789974428059 0.7765171413166 0.7698864925325 0.7924280641870 0.789628259467 0.7732159684544
MRNN 0.8012121900939 0.7994518346619 0.7939973744164 0.828283355107 0.8303854640068 0.823715692404
CSAI 0.86464530892 0.85919145690 0.83722349351 0.837759164790 0.8400507785337 0.8317807876933
[…]
@@
Evaluating Non-uniform Masking The effect of different configurations on model performance was evaluated by non-uniformly masking various combinations of the training, validation, and test sets, while examining the number of training epochs required to achieve the reported performance for each. The results for the BRITS and BRITS-GRU baseline models are shown in Table 3. These results demonstrate that consistently applying non-uniform masking across all data partitions yielded the best performance. This suggests that the masking strategy effectively adjusts the representation of different features to optimally leverage the data’s distribution, enhancing the model’s ability to handle the inherent heterogeneity of the dataset.
Table 3: Performance comparison across masking permutations showing the impact of different non-uniform masking strategies on both imputation and classification, using the BRITS and BRITS-GRU models. Train only: Non-uniform masking applied solely to the training set. All: non-uniform masking is applied across all training, validation and test sets. None: a baseline with no non-uniform masking. Val only: Non-uniform masking is applied only to the validation set. Val Test: Non-uniform masking is applied to both the validation and test sets. Test only: Non-uniform masking is applied exclusively to the test set. Under the ’All’ configuration, non-uniform masking is applied during training, validation, and testing consistently demonstrates superior performance (bold font).

### v4

@@ IV-B Non-Uniform Masking Strategy
  • $\boldsymbol{R_{\text{factor}}}(d|U,I)$ is the adjustment factor for feature $d$, conditioned by $U$ and $I$.
  • $\boldsymbol{P_{\text{dist}}}(d)$ is $d$’s missingness propability distribution.
The overall masking proportion is then adjusted to ensure consistency with the uniform masking rate $U$, while retaining the non-uniform characteristics of the individual features. The pseudocode for this algorithm is given below, producing a masked matrix $\boldsymbol{X_{M}}$.
  Input: Dataset $\boldsymbol{X}$ with $D$ features, masking rate $U$, weighting parameter $I$
@@
  for each feature $d$ in $D$ do
     $\boldsymbol{P_{\text{dist}}}(d)\leftarrow\text{compute}(\boldsymbol{x^{d}})$
@@
     $\boldsymbol{P_{\text{nu}}}(d)\leftarrow\boldsymbol{R_{\text{factor}}}(d)\times\boldsymbol{P_{\text{dist}}}(d)$
@@
  $U\leftarrow f(\boldsymbol{P_{\text{nu}}},\boldsymbol{X})$
@@
Algorithm 1 non-uniform-mask
IV-C Learning
CSAI is trained in an unsupervised manner by stochastically masking non-missing values using our non-uniform masking strategy and learning to impute them. At evaluation time, the trained RNN can impute both real and simulated missing values. CSAI’s training iterates over mini-batches of the input data, where each mini-batch comprises a set of samples with $T$ time steps. Our imputation loss function $\mathcal{L}_{imp}$ maintains the objective of minimising the reconstruction error $\ell_{\text{obs}}$ between the imputed and ground truth vectors $\boldsymbol{x}$ and $\boldsymbol{\hat{x}}$, while maintaining consistency loss $\ell_{\text{con}}$ between the forward and backward imputed estimates $\overrightarrow{\hat{\boldsymbol{x}}}$ and $\overleftarrow{\hat{\boldsymbol{x}}}$ of our bi-directional RNN.
CSAI’s training iterates over mini-batches of the input data, where each mini-batch comprises a set of samples with $T$ time steps. Our imputation loss function $\mathcal{L}_{imp}$ maintains the objective of minimising the reconstruction error $\ell_{\text{obs}}$ between the imputed and ground truth vectors $\boldsymbol{x}$ and $\boldsymbol{\hat{x}}$, while maintaining consistency loss $\ell_{\text{con}}$ between the forward and backward imputed estimates $\overrightarrow{\hat{\boldsymbol{x}}}$ and $\overleftarrow{\hat{\boldsymbol{x}}}$ of our bi-directional RNN as follows:
$\mathcal{L}_{imp}=\frac{1}{|\mathcal{B}|}\sum_{i\in\mathcal{B}}\sum_{t\in T_{i}}\lambda\Big{[}$ $\boldsymbol{M}_{it}\odot\ell_{\text{obs}}(\boldsymbol{x}_{it},\hat{\boldsymbol{x}}_{it})$
$+(1-\boldsymbol{M}_{it})\odot\ell_{\text{con}}(\overrightarrow{\hat{\boldsymbol{x}}_{it}},\overleftarrow{\hat{\boldsymbol{x}}_{it}})\Big{]}$ (23)
Where $\mathcal{B}$ is the mini-batch size, $T$ is the set of time steps over which the loss function is applied, $\boldsymbol{M}$ is the masking matrix and $\lambda$ is a hyperparameter that balances the two loss terms. Both $\ell_{obs}$ and $\ell_{cons}$ use the mean absolute error (MAE) to measure the difference between the components of each loss function.
Since CSAI can be used as an end-to-end pipeline to perform imputation and prediction, we further define prediction loss using binary cross-entropy as follows:
$\mathcal{L}_{\text{pred}}=-\sum_{n=1}^{N}y^{(n)}\log(\hat{y}^{(n)})$ (24)
The resulting combined loss is computed as:
$\mathcal{L}_{\text{combined}}=\alpha\mathcal{L}_{\text{imp}}+\beta\mathcal{L}_{\text{pred}}$ (25)
Where $\beta$ With $\alpha$ and $\beta$ are preset parameters representing the weight of the respective loss component on $\mathcal{L}_{\text{combined}}$. Using CSAI solely as imputer is done by setting $\beta$ to zero. With this loss function, CSAI uses backpropagation in an end-to-end manner to optimise model parameters during training.
V Experimental Evaluation
We conducted several experiments to evaluate the proposed methodologies and their impact on CSAI’s performance on widely-used healthcare benchmark datasets. First, we describe the datasets, followed by the experimental design and implementation details, and conclude with the results.
V-A Datasets
We used three healthcare benchmarks obtained from three widely used databases. Each of the datasets used has a different data distribution and characteristics as shown in Table I.
Dataset Size Avg Baseline Missingness Feature Correlation (Static; Categorical) Variables
eICU 30,680 x 20 40.53% 0.14 (6; 3)
MIMIC_59 21,128 x 59 61% 0.17 (0;0)
Physionet 3,997 x 35 51% 0.12 (0; 0)
TABLE I: Characteristics of the four datasets: Size is described as number of samples $\times$ number of variables. Average baseline missingness of all features. Feature correlation is the average feature-wise Pearson correlation coefficient. (static; categorical) the number of static (as opposed to dynamic) and categorical (as opposed to numerical) variables in the dataset.
  1. wide
    eICU [^1] https://physionet.org/content/eicu-crd/2.0/ [27] is a multicentre database with anonymised patient records from intensive care units (ICU) obtained from more than 200 hospitals. eICU is publicly available after registration. We followed the only benchmark extract available for eICU [28].
  2. wiide
@@ V-A Datasets
    MIMIC-III [^2] https://physionet.org/content/mimiciii/1.4/ [29] is an extensive, single-hospital database of over 40,000 critical care patients. We used MIMIC-III to increase the heterogeneity of our datasets with different features types and data dimensionality and followed a well-cited benchmark [30] to produce a dataset comprising 59 numerical variables for 21,128 samples.
  3. wiiide
@@
    PhysioNet Challenge 2012 dataset is a public medical benchmarking dataset provided by [31], containing records of 4000 48-hour hospital stays, allowing unbiased evaluations of different model performances. This benchmark comprises 35 numerical variables for 3,997 samples.
For all data sets, we reproduced the available benchmarking, skipping the steps that remove all-NAN samples to retain the data with its original missingness.
V-B Experimental Design
We conducted two sets of experiments to test the methodologies proposed in this work. The first comprises benchmark comparison experiments comparing CSAI’s imputation and subsequent classification performance against the current state of the art on two scales:
  1. wide
    Experiment I: compares CSAI with the best-performing RNN models using the three benchmark datasets and three masking rations (5%, 10% and 20%). Here, CSAI is compared with our backbone bidirectional RNN model BRITS [14], the Gated Recurrent Unit with Decay (GRU-D) [13], the Variational Recurrent Imputation Network (V-RIN) [23], and the Multi-directional Recurrent Neural Network (M-RNN) model [15]. In this experiment, two tasks were undertaken to ensure a fair comparison: 1) because the original BRITS implementation uses LSTM cells while all other baseline models use GRU cells, we implemented a GRU-based BRITS model to ensure a balanced evaluation across similar architectures, including its output along with the original BRITS (BRITS-GRU in the results), 2) we employed our nonuniform masking algorithm as a pre-processing step for all models to avoid any bias arising from the masking step.
  2. wiide
    Experiment II: is a large-scale experimental evaluation with different flavours of neural imputation models using the Physionet dataset with a 10% masking ratio. This experiment was conducted using the PyPOTS library [^3] https://pypots.com/about/ [32], an open-source Python toolkit for benchmarking and analysing incomplete multivariate time series. PyPOTS provides access to 29 imputation algorithms, including CSAI and a number of benchmark datasets, including Physionet. In this experiment, CSAI is compared against a broad range of neural network imputation models, such as Transformers, CNNs, GNNs, and diffusion models, as well as RNNs and standard statistical imputation methods. As non-uniform masking is not yet integrated into PyPOTS, we rely on the available masking setup, aiming to evaluate the contribution of CSAI’s temporal decay and attention mechanism on the resulting performance.
    The reader should note that although a comparison with generative models through $E^{2}$ [33] would have provided valuable insight, particularly given its lack of its quantitative comparison with BRITS, the publicly available implementation of $E^{2}$ is incompatible with our current setup and is also not part of PyPOTS.
The second set of experiments are ablation studies performed using the public Physionet dataset. First, we establish the effect of non-uniform masking across different data partitions, namely the training, validation, and test sets. The aim is to establish whether this approach actually improves a model’s ability to generalise and handle missing data in a more realistic and robust manner. Finally, we evaluate the effect of the weighting parameter, which controls the degree of emphasis placed on different features in the non-uniform masking strategy. This parameter allows for fine-tuning of the feature representation during model training, directly influencing how the model handles missing data and we evaluate the effect on CSAI’s imputation and classification performance under different weighting parameter values.
V-C Experimental Setup
Environment: We trained our models on an HPC node equipped with an NVIDIA A100 40GB GPU, running Ubuntu 20.04.6 LTS (Focal Fossa). The experiments were carried out using Python 3.8.16. For the first set of experiments comparing CSAI with other RNN models, we directly used the source code for BRITS, GRU-D, and MRNN, obtained from their respective repositories, along with CSAI. For the broader set of experiments comparing CSAI with all neural imputation models, we utilised the PyPOTS implementation of all models, running all experiments through the PyPOTS package installed on our HPC node. Full package details can be found in our repository. To help reproduce our Python environment easily, we freeze the development environment with Anaconda and save it into file for reference, which is available in the project’s GitHub repository. To ensure reproducibility of our results, we release all data preprocessing scripts, model implementations, and hyperparameter search configurations.
Dataset Pre-processing and Missingness Patterns Simulation: Each dataset was normalised to zero mean and unit variance. For each of the four datasets, we created three versions by further masking 5%, 10% and 20% cells in addition to the missingness already present in the original datasets. These masked cells have known ground truths and will form the basis of our comparison of imputation performance. To more effectively simulate EHR missingness scenarios, our first set of experiments use our non-uniform masking as a pre-processing step to enable capturing patterns representative of common missingness data mechanisms in EHRs to the masked cells. In our wider experiments using PyPOTs, we used the package’s PyGrinder toolbox  [^4] https://pypots.com/ecosystem/#PyGrinder, which implements structured missingness patterns that reflect the dynamic nature of time series data [34, 35, 36].
Hyper-parameter Optimisation: The performance of deep-learning models highly depends on the settings of hyper-parameters. To make fair comparisons across imputation methods and draw impartial conclusions, hyper-parameter optimisation is applied to all imputation algorithms using PyPOTS.
Training: Except the PhysioNet dataset which already contains separated time series samples, all other datasets are split into training, validation, and test sets. Randomly, we selected 10% of each dataset for validation and another 10% for testing, training the models on the remaining data. We used the Adam optimiser and set the number of RNN hidden units to 108 for all models. The batch size is 64 for PhysioNet and 128 for the other datasets. A 5-fold cross-validation method was implemented to evaluate the models.
Downstream Task Design: To evaluate the impact of imputation on downstream analysis, we further perform a classification task to predict in-hospital mortality. In the Physionet dataset, each sample has a label indicating whether the patient is deceased in the ICU. For all other datasets, we extracted the results for each patient according to the code provided in the benchmark. To demonstrate the flexibility of the model and its ability to perform using different classification architectures, we used different methodologies to implement the classifiers in our experiments. In Experiment I, classification was performed in an end-to-end manner by adding a classification layer to each architecture, utilising the hidden states from the imputation network to feed into the classification layer. In Experiment II, the imputation models were used to generate complete datasets, which were subsequently fed into three different classifiers for comparison: an XGBoost, an RNN and a Transformer classifier.
Evaluation Metrics: To assess imputation performance, we use mean absolute error (MAE) [14] defined as:
$\text{MAE}=\frac{1}{n}\sum_{i=1}^{n}\left|y_{i}-\hat{y}_{i}\right|$
Where:
  • $n$ is the number of data points,
  • $y_{i}$ is the actual value for the $i$-th data point,
  • $\hat{y}_{i}$ is the predicted value for the $i$-th data point.
For classification, we use the Area Under the Receiver-Operator Curve (ROC-AUC) to report performance.
@@ V-C Experimental Setup
TABLE II: Imputation performance using 5%, 10%, and 20% masking ratios on three datasets. The model with the lowest MAE in each setup is highlighted in bold.
eICU (MAE) MIMIC_59 (MAE) PhysioNet (MAE)
5% 10% 20% 5% 10% 20% 5% 10% 20%
V-RIN 0.24161 ± 0.015 0.24254 ± 0.013 0.25214 ± 0.019 0.15457 ± 0.007 0.13818 ± 0.017 0.33697 ± 0.010 0.26163 ± 0.015 0.27372 ± 0.010 0.29997 ± 0.018
BRITS 0.16699 ± 0.014 0.17053 ± 0.020 0.17681 ± 0.009 0.15195 ± 0.018 0.14023 ± 0.009 0.34039 ± 0.019 0.25634 ± 0.013 0.26762 ± 0.017 0.28722 ± 0.014
BRITS_GRU 0.17232 ± 0.010 0.17124 ± 0.019 0.17691 ± 0.013 0.14793 ± 0.016 0.14193 ± 0.015 0.34169 ± 0.017 0.25129 ± 0.012 0.26216 ± 0.011 0.28288 ± 0.018
GRUD 0.22274 ± 0.018 0.22560 ± 0.010 0.23098 ± 0.020 0.30447 ± 0.012 0.28704 ± 0.014 0.48670 ± 0.017 0.49406 ± 0.015 0.49779 ± 0.020 0.50952 ± 0.018
MRNN 0.47036 ± 0.015 0.47998 ± 0.017 0.50065 ± 0.020 0.30573 ± 0.013 0.28342 ± 0.012 0.47198 ± 0.015 0.54671 ± 0.013 0.55647 ± 0.014 0.57230 ± 0.017
CSAI 0.15967 ± 0.017 0.16149 ± 0.011 0.16637 ± 0.015 0.13119 ± 0.009 0.11291 ± 0.008 0.30976 ± 0.014 0.24602 ± 0.014 0.25747 ± 0.017 0.27476 ± 0.019
V-D Experimental Results
Experiment I: Comparison with State-of-the-Art RNN Models on Healthcare Benchmarks: Our evaluation of CSAI against four state-of-the-art RNN imputation models on all three healthcare benchmarks across three masking ratios is shown in Tables II and III. Table II shows the imputation performance, where CSAI consistently outperforms other models in all data sets and masking ratios (5%, 10%, and 20%). The best performance for all models is observed on the eICU and MIMIC_59 datasets, which, despite high missingness rates, offers a large number of samples for training. The Physionet dataset, while having the lowest baseline missingness rate, provides significantly fewer samples compared to the other datasets, limiting model performance across the board. The table also shows that the performance gap between CSAI and other models widens as the masking ratio increases, leading to conditions of high data loss. This is particularly pronounced in the highly-dimensional dataset MIMIC_59 at the 20% missingness ratio. For V-RIN, BRITS, and BRITS_GRU, the MAE is relatively stable across masking ratios but remains consistently higher than CSAI. GRUD and MRNN show notably higher MAE values, especially at higher masking ratios.
Table III further demonstrates CSAI’s strong classification performance, achieving the highest AUC scores across various masking ratios in the eICU, PhysioNet, and MIMIC_59 datasets. The difference between CSAI’s AUCs and those of other models is highest under the Physionet dataset, which has the smallest number of samples. Across all three datasets, CSAI’s AUC only slightly decreases as the masking ratio increases to 20%. Although a similar stability is observed in BRITS and V-RIN, the AUCs are consistently lower than CSAI’s.
@@ V-D Experimental Results
TABLE III: Classification performance using 5%, 10%, and 20% masking ratios on three datasets. The model with the highest AUC in each setup is highlighted in bold.
eICU (AUC) MIMIC_59 (AUC) PhysioNet (AUC)
5% 10% 20% 5% 10% 20% 5% 10% 20%
V-RIN 0.88766 ± 0.012 0.88422 ± 0.013 0.88461 ± 0.015 0.83280 ± 0.010 0.83312 ± 0.012 0.82733 ± 0.014 0.83433 ± 0.011 0.82916 ± 0.010 0.82548 ± 0.015
BRITS 0.88669 ± 0.011 0.88521 ± 0.013 0.88569 ± 0.012 0.82821 ± 0.010 0.82779 ± 0.012 0.82414 ± 0.013 0.82214 ± 0.014 0.81175 ± 0.012 0.82181 ± 0.015
BRITS_GRU 0.88944 ± 0.012 0.88864 ± 0.015 0.88605 ± 0.011 0.83188 ± 0.010 0.83074 ± 0.012 0.82688 ± 0.013 0.80677 ± 0.014 0.81926 ± 0.013 0.80651 ± 0.015
GRUD 0.86495 ± 0.013 0.86456 ± 0.014 0.85801 ± 0.011 0.82768 ± 0.012 0.82640 ± 0.014 0.82304 ± 0.012 0.78997 ± 0.011 0.77652 ± 0.013 0.76989 ± 0.015
MRNN 0.87786 ± 0.011 0.87626 ± 0.014 0.87342 ± 0.015 0.82106 ± 0.012 0.81708 ± 0.011 0.81282 ± 0.013 0.80121 ± 0.014 0.79945 ± 0.013 0.79400 ± 0.015
CSAI 0.88952 ± 0.012 0.88977 ± 0.011 0.88795 ± 0.015 0.83524 ± 0.014 0.83367 ± 0.013 0.83109 ± 0.012 0.86465 ± 0.014 0.85919 ± 0.013 0.83722 ± 0.015
Experiment II: Large-scale Comparison Using PyPOTs: Table IV shows classification results for 24 neural imputation models and four statistical imputers. This experiment uses the imputed data generated by each model as input to a classifier. We show the results for three classifiers, and XGBoost, RNN and a Transformer classifier. The reader should not that non-uniform masking was not implemented into CSAI here, since it is yet to be integrated into a callable capability in the PyPOTs library and it is therefore not possible to use it with any other imputation models apart from CSAI.
Overall, the XGBoost classifier achieved the best performance, with a mean AUC of 0.823 compared to 0.660 for the RNN classifier and 0.667 for the Transformer. With XGBoost, using CSAI as an imputer produced the best classification performance. Since XGBoost is a non-temporal model, it has clearly benefited from CSAI’s ability to encode sequential dependencies (across features and time) as informative features that complement XGBoost’s strengths in feature-based learning. For both the RNN and Transformer classifiers, diffusion models produced the highest AUCs, largely due to their ability to offer RNN and Transformer imputers the advantage of smoother interpolations that likely enhance feature consistency across time steps, which is crucial for sequential models. Despite this, CSAI ranked highly under these classifiers, producing competitive AUC performance as a lightweight, efficient model that avoids the computational burden of diffusion models.
TABLE IV: Classification performance on the Physionet dataset using 10% masking ratio. The imputation was performed using the PyPOTS implementation of the major Transformer, RNN, CNN, GNN, diffusion, and statistical models. For comparison, imputation using XGBoost’s internal imputer resulted in ROC_AUC 0.771 $\pm$ 0.000
Model Type Model XGB Classifier RNN Classifier Transformer Classifier
Transformers iTransformer 0.852 $\pm$ 0.000 0.692 $\pm$ 0.073 0.685 $\pm$ 0.032
SAITS 0.851 $\pm$ 0.000 0.658 $\pm$ 0.076 0.668 $\pm$ 0.037
ETSformer 0.820 $\pm$ 0.000 0.734 $\pm$ 0.030 0.678 $\pm$ 0.013
PatchTST 0.848 $\pm$ 0.000 0.703 $\pm$ 0.064 0.698 $\pm$ 0.061
Crossformer 0.836 $\pm$ 0.000 0.683 $\pm$ 0.051 0.644 $\pm$ 0.026
Informer 0.836 $\pm$ 0.000 0.701 $\pm$ 0.021 0.653 $\pm$ 0.029
Autoformer 0.768 $\pm$ 0.000 0.597 $\pm$ 0.009 0.625 $\pm$ 0.008
Pyraformer 0.829 $\pm$ 0.000 0.739 $\pm$ 0.026 0.654 $\pm$ 0.047
Transformer 0.833 $\pm$ 0.000 0.723 $\pm$ 0.036 0.626 $\pm$ 0.036
RNN BRITS 0.841 $\pm$ 0.000 0.667 $\pm$ 0.066 0.646 $\pm$ 0.074
MRNN 0.760 $\pm$ 0.000 0.611 $\pm$ 0.008 0.619 $\pm$ 0.015
GRUD 0.840 $\pm$ 0.000 0.701 $\pm$ 0.075 0.704 $\pm$ 0.035
CSAI 0.860 $\pm$ 0.000 0.702 $\pm$ 0.044 0.704 $\pm$ 0.023
CNN TimesNet 0.823 $\pm$ 0.000 0.687 $\pm$ 0.087 0.710 $\pm$ 0.043
MICN 0.824 $\pm$ 0.000 0.700 $\pm$ 0.071 0.686 $\pm$ 0.050
SCINet 0.818 $\pm$ 0.000 0.663 $\pm$ 0.055 0.720 $\pm$ 0.034
GNNs StemGNN 0.809 $\pm$ 0.000 0.534 $\pm$ 0.124 0.701 $\pm$ 0.025
FreTS 0.840 $\pm$ 0.000 0.693 $\pm$ 0.035 0.674 $\pm$ 0.018
Koopa 0.835 $\pm$ 0.000 0.509 $\pm$ 0.113 0.628 $\pm$ 0.036
DLinear 0.851 $\pm$ 0.000 0.689 $\pm$ 0.038 0.611 $\pm$ 0.053
FiLM 0.825 $\pm$ 0.000 0.565 $\pm$ 0.095 0.706 $\pm$ 0.046
Diffus. CSDI 0.853 $\pm$ 0.000 0.553 $\pm$ 0.108 0.762 $\pm$ 0.019
US-GAN 0.839 $\pm$ 0.000 0.708 $\pm$ 0.019 0.610 $\pm$ 0.043
GP-VAE 0.816 $\pm$ 0.000 0.745 $\pm$ 0.056 0.701 $\pm$ 0.041
Stats Mean 0.763 $\pm$ 0.000 0.620 $\pm$ 0.017 0.598 $\pm$ 0.025
Median 0.772 $\pm$ 0.010 0.619 $\pm$ 0.013 0.605 $\pm$ 0.025
LOCF 0.795 $\pm$ 0.033 0.634 $\pm$ 0.074 0.644 $\pm$ 0.061
Linear 0.807 $\pm$ 0.036 0.640 $\pm$ 0.076 0.669 $\pm$ 0.069
Ablation Study Figure 4 demonstrates a clear progression in CSAI’s performance as its components are incrementally added. Starting from the baseline, the model’s MAE improves with each addition, highlighting the impact of the domain-informed temporal decay, attention-based initialisation mechanism, and non-uniform masking. The final configuration, with all components included, achieves the lowest MAE across all three datasets. These results indicate that each component plays a distinct role in enhancing imputation accuracy, with the full model yielding the best performance.
Figure 4: Visualisation of the incremental increase in imputation performance (decrease in MAE) as the different components of CSAI are incrementally added. The figure shows CSAI’s MAE using the three datasets and 10% masking ratio and the following variations: a) baseline model (i.e. a BRITS architecture), b) adding the temporal decay function, c) adding the attention-base initialisation mechanism, and finally d) using non-uniform masking.
@@
We further evaluate the effect of non-uniform masking on algorithm performance by using the masking strategy on different combinations of the training, validation and test sets and examining the number of training epochs required to achieve the reported performance for each. For objective evaluation of our non-uniform masking strategy, we conducted all experiments using the BRITS and BRITS-GRU baseline models, with the results shown in Table V. These results demonstrate that consistently applying non-uniform masking across all data partitions (training, validation and testing) yields the best performance, suggesting that the masking strategy effectively adjusts the representation of different features to optimally leverage the data distribution, enhancing the model’s ability to handle the inherent heterogeneity of the dataset.
TABLE V: Performance comparison of applying our non-uniform masking strategy on different combinations of the dataset: training, validation and testing. Under the ’All’ configuration, non-uniform masking is applied during training, validation, and testing consistently reports the performance (bold font).